4.2+(3.8 x 3.8 x 1 over 4) - 5 =


Answer by Alan3354(21584)

4.2+(3.8 x 3.8 x 1 over 4) - 5 =
=4.2 +(14.44*(1/4)) - 5
=4.2 + 3.61 - 5
= 7.81 - 5
= 2.81